OpenStudy (anonymous):

If the substitution u=(x/2) is made, the integral...

OpenStudy (anonymous):

\[\int\limits_{2}^{4}(\frac{ 1-\frac{ x }{ 2 }^2) }{ x }dx=? \]

OpenStudy (anonymous):

I have to clarify, the \[\frac{ x ^{2} }{ 2 } \] should be \[(\frac{ x }{ 2 })^2\]

OpenStudy (anonymous):

\[\large\int_2^4\frac{1-\left(\frac{x}{2}\right)^2}{x}dx\;\;?\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[u=\frac{x}{2}\Rightarrow x=2u\\ du=\frac{1}{2}dx\\ 2\;du=dx\] \[\int_1^2\frac{1-u^2}{2u}(2\;du)=\int_1^2\frac{1-u^2}{u}\;du\]

OpenStudy (anonymous):

If you were wondering why the limits change: \[\text{lower: }x=2\;\Rightarrow\;u=\frac{2}{2}=1\\ \text{upper: }x=4\;\Rightarrow\;u=\frac{4}{2}=2\]

OpenStudy (anonymous):

Ah of course! I forgot I could express x in terms of u. Thank you so much!

OpenStudy (anonymous):

You're welcome!